Evaluate the following functions in differential operator form.

  1. D(x2+2x-3) Ans. (2x+2)
  2. D2(xe3x-e4x) Ans. 9xe3x+6e3x-16e4x
1
Expert's answer
2022-01-18T06:22:19-0500

"D(x^2 + 2x + 3) \\
\text{Since}\, D = \dfrac{d}{dx} \\
\text{So, we have,}\\
\dfrac{d}{dx} (x^2 + 2x + 3) \\
= \dfrac{d}{dx} (x^2) + \dfrac{d}{dx} (2x) 
+ \dfrac{d}{dx} (3) \\
= 2x + 2 \\

\\

D^2(xe^{3x} - e^{4x})\\
\text{Since} \, D = \dfrac{d}{dx} \\
\text{So, we have,}\\
\dfrac{d^2}{dx^2} (xe^{3x} - e^{4x})\\
\dfrac{d}{dx} (\dfrac{d}{dx} (xe^{3x} - e^{4x}))\\
= \dfrac{d}{dx} (\dfrac{d}{dx} xe^{3x}
- \dfrac{d}{dx} e^{4x})\\
\text{Recall the product rule of differentiation}\\
UV = Vdu + Udv\\
\text{By comparison let} \, u = x, v = e^{3x}\\
\text{Upon differentiation. We have,}\\
\dfrac{d}{dx} (e^{3x} + 3xe^{3x} - 4e^{4x})\\
= \dfrac{d}{dx} (e^{3x}) + 3 \dfrac{d}{dx} (xe^{3x})
- 4 \dfrac{d}{dx} (e^{4x})\\
= 3e^{3x} + 3(e^{3x} + 3xe^{3x}) - 4(4e^{4x})\\
= 3e^{3x} + 3e^{3x} + 9xe^{3x} - 16e^{4x}\\
= 9xe^{3x} + 6e^{3x} - 16e^{4x}"